# Selective Fine-tuning on LLM-labeled Data May Reduce Reliance on Human Annotation: A Case Study Using Schedule-of-Event Table Detection

### v1

@@ 4.2 Fine-tuned Models
We conduct three sets of fine-tuning experiments using PaLM-2 models. We first conduct fine-tuning on PaLM-2 model using the non-expert human annotation obtained as described in section 3.3. This serves as a strong benchmark for our second and third fine-tuning experiments that use gemini-pro 1.0 based annotations for fine-tuning PaLM-2. The second and third finetuning experiments differ in the sense that while one of the experiments use entirety of gemini-pro’s labels during fine-tuning, the other experiment is fine-tuned on only those gemini-pro labels where there is a consensus between Gemini-pro’s inference for JSON and text-based representation of a given table.
As seen in table 2, fine-tuning PaLM-2 with labels generated by Gemini-pro 1.0 leads to improvements over the base PaLM-2 model’s performance. However, this improvement is nuanced. When the model is fine-tuned using the entirety of gemini-pro’s labels, it improves over the baseline PaLM-2 model. However, the precision, f1-score, and accuracy compared to the standalone gemini-pro 1.0 model remains inferior. Optimal fine-tuning is achieved through only incorporating labels for which there is a consensus between gemini-pro’s JSON and text-based inferences for a table. This fine-tuned variant (table 2 last row), not only exceeds the performance of the base PaLM-2 and gemini-pro 1.0 models, but also narrows the gap to the precision and f1-score of PaLM-2 fine-tuned with human labels, though it doesn’t completely bridge it. Furthermore, our fine-tuned models surpass the performance of naive ensemble approaches combining gemini-pro 1.0 and PaLM-2, as detailed in Appendix C. Yet, it is worth noting that the naive ensembles still offer better results than either standalone base model. These findings underscore the value of naive ensembles for preliminary analysis and for scenarios where fine-tuning isn’t feasible, while also highlighting the potentially superior results attainable with fine-tuned models.
@@
Figure 1: Precision and F1 Score Across Models for 91 protocols in test set. Bubble sizes represent the number of SoE tables within a protocol. PaLM-2 models fine-tuned with human labels and consensus-based gemini-pro 1.0 labels achieve a median precision of 100% and F1 score of 1.
@@ 5 Discussion
Our study proposes a novel approach to fine-tuning Large Language Models for specialized domains where labels can be especially difficult to obtain. The approach of utilizing noisy labels from an LLM (gemini-pro) for fine-tuning another LLM (PaLM-2) demonstrates a scalable and cost-effective alternative to conventional expert annotation processes. Notably, the introduction of a label filtering mechanism, which selects labels for fine-tuning only when there is agreement between dual data representations (JSON and text) of a table, leads to substantial improvements in model performance. This method not only outperforms the base model and the label-generating LLM, it also approaches the performance level of models fine-tuned with human annotations on our table classification task, highlighting the potential of our fine-tuning strategy to effectively leverage auto-generated labels. While we use JSON and text based consensus approach as a proxy for selecting high quality training labels for our specific table classification task, any multi-modal data representation can serve a similar purpose in filtering out potentially noisy labels.
@@
Despite its strengths, our approach has several limitations. The task of SoE table detection is highly specialized and relies on data from our internal Clinical Trial Management System (CTMS) software, which may not fully capture the variety and complexity of clinical trial protocols encountered in broader applications. This specificity could limit the generalizability of our findings to other types of documents or on different table classification tasks. If LLMs perform poorly across the board on a specific task, automated generation of labels may not be feasible even with powerful models like gemini-pro. Moreover, while our study underscores the feasibility and effectiveness of using LLM-generated labels for fine-tuning, it lacks a direct comparison with a baseline model fine-tuned on expert annotations due to the high costs and resource requirements associated with obtaining such annotations. This comparison could have provided a clearer benchmark for evaluating the relative performance of our approach. Importantly, the reliance on auto-label generation and consensus-based fine-tuning may introduce or perpetuate biases inherent in the models used for label generation. Depending on specific context and fine-tuning task, these biases may manifest as demographic, entity, or domain-specific biases, affecting the accuracy and fairness of the fine-tuned model, particularly in sensitive domains like healthcare. Thus, for settings where there is a concern regarding bias and fairness in model generated labels and outputs, comprehensive fairness and bias evaluation specific to auto-labeling and fine-tuning steps are essential for detecting and mitigating biases in auto-generated labels and fine-tuned models. This can potentially mitigate such concerns for auto-generated labels and model fine-tuned on such labels paving the way for broader adoption of LLMs in data-rich, but expert-scarce domain like healthcare, where processes often rely heavily on manual workflows.
@@ Appendix A Schedule of Event Tables
We provide a sample SoE table (refer table 3) based on NIH template and two non-SoE tables (refer tables 4 and 5) below. The first table has clear screening, treatment and follow-up period. It specifies various visit with time information as well as window during which the visit can take place. The second table looks like an SoE table in terms of structure, but it doesn’t have clearly demarcated screening, treatment and follow-up. Further, it only specifies specific lab tests at the start of the diagnosis and completion of therapy and lacks treatment period information. Often this table would require protocol digitization specialists to look at additional context (like surrounding texts on the page) in the protocol to determine whether or not it is a SoE table. The last table specifies pharmacokinetic collections and is not a SoE table (see prompts in appendix B which we wrote in consultation with the digitizers for SoE tables)

### v2

@@ 4.2 Fine-tuned Models
We conduct four sets of fine-tuning experiments using PaLM-2 models. We first conduct fine-tuning on PaLM-2 model using the non-expert human annotation obtained as described in section 3.3. This serves as a strong benchmark for our remaining fine-tuning experiments that use gemini-pro 1.0 based annotations for fine-tuning PaLM-2. The second and third fine-tuning experiments differ in the sense that while one of the experiments use entirety of gemini-pro’s labels during fine-tuning, the other experiment is fine-tuned on only those gemini-pro labels where there is a consensus between gemini-pro’s inference for JSON and text-based representation of a given table. Finally, we explore a hybrid approach to further refine the fine-tuning process. Instead of discarding the samples lacking consensus between JSON and text-based inferences, we replace the Gemini Pro labels for those samples with human annotations. This results in a fine-tuning dataset consisting of 2518 Gemini-annotated samples and 282 human-annotated ones.
As seen in table 3, fine-tuning PaLM-2 with labels generated by Gemini-pro 1.0 leads to improvements over the base PaLM-2 model’s performance. However, this improvement is nuanced. When the model is fine-tuned using the entirety of gemini-pro’s labels, it improves over the baseline PaLM-2 model. However, the precision, f1-score, and accuracy compared to the standalone gemini-pro 1.0 model remains inferior. Optimal fine-tuning is achieved through only incorporating labels for which there is a consensus between gemini-pro’s JSON and text-based inferences for a table. This fine-tuned variant (table 3 second last row), not only exceeds the performance of the base PaLM-2 and gemini-pro 1.0 models, but also narrows the gap to the precision and f1-score of PaLM-2 fine-tuned with human labels. Our hybrid labeling strategy (table 3 last row), incorporating human annotations for cases lacking consensus, further bridges the gap with the model fine-tuned on human annotations.
Finally, our fine-tuned models surpass the performance of naive ensemble approaches combining gemini-pro 1.0 and PaLM-2, as detailed in Appendix C. Yet, it is worth noting that the naive ensembles still offer better results than either standalone base model. These findings underscore the value of naive ensembles for preliminary analysis and for scenarios where fine-tuning isn’t feasible, while also highlighting the potentially superior results attainable with fine-tuned models.
@@
Figure 2: Precision and F1 Score across models for 91 protocols in test set (hybrid approach not shown for brevity.) Bubble sizes represent the number of SoE tables within a protocol. PaLM-2 models fine-tuned with human labels and consensus-based gemini-pro 1.0 labels achieve a median precision of 100% and F1 score of 1.
@@ 5 Discussion
Our study proposes a novel approach to fine-tuning Large Language Models for specialized domains where labels can be especially difficult to obtain. The approach of utilizing noisy labels from an LLM (gemini-pro) for fine-tuning another LLM (PaLM-2) demonstrates a scalable and cost-effective alternative to conventional expert annotation processes. Notably, the introduction of a label filtering mechanism, which selects labels for fine-tuning only when there is agreement between dual data representations (JSON and text) of a table, leads to substantial improvements in model performance. This method not only outperforms the base model and the label-generating LLM, it also approaches the performance level of models fine-tuned with human annotations on our table classification task, highlighting the potential of our fine-tuning strategy to effectively leverage auto-generated labels. The filtering mechanism also provides an easy way to identify potential samples that may require expert labeling. As demonstrated by our hybrid labeling approach, this can significantly reduce the annotation workload while potentially bringing fine-tuned models on par with models fine-tuned on human labels. This targeted approach to human annotation allows for a more efficient allocation of resources, maximizing the benefits of both automated and expert labeling. While we use JSON and text based consensus approach as a proxy for selecting high quality training labels for our specific table classification task, any multi-modal data representation can serve a similar purpose in filtering out potentially noisy labels.
@@
Despite its strengths, our approach has several limitations. The task of SoE table detection is highly specialized and relies on data from our internal Clinical Trial Management System (CTMS) software, which may not fully capture the variety and complexity of clinical trial protocols encountered in broader applications. This specificity could limit the generalizability of our findings to other types of documents or on different table classification tasks. If LLMs perform poorly across the board on a specific task, automated generation of labels may not be feasible even with powerful models like gemini-pro. Moreover, while our study underscores the feasibility and effectiveness of using LLM-generated labels for fine-tuning, it lacks a direct comparison with a baseline model fine-tuned on expert annotations due to the high costs and resource requirements associated with obtaining such annotations. This comparison could have provided a clearer benchmark for evaluating the relative performance of our approach. Importantly, the reliance on auto-label generation and consensus-based fine-tuning may introduce or perpetuate biases inherent in the models used for label generation. Depending on specific context and fine-tuning task, these biases may manifest as demographic, entity, or domain-specific biases, affecting the accuracy and fairness of the fine-tuned model, particularly in sensitive domains like healthcare. While our hybrid approach, incorporating human annotations for challenging or low-confidence cases, can help mitigate some of these biases, comprehensive fairness and bias evaluation specific to both the auto-labeling and fine-tuning steps are still essential for detecting and mitigating biases in auto-generated labels and fine-tuned models. This can pave the way for broader adoption of LLMs in data-rich, but expert-scarce domains like healthcare, where processes often rely heavily on manual workflows.
@@ Appendix A Schedule of Event Tables
We provide a sample SoE table (refer table 4) based on NIH template and two non-SoE tables (refer tables 5 and 6) below. The first table has clear screening, treatment and follow-up period. It specifies various visit with time information as well as window during which the visit can take place. The second table looks like an SoE table in terms of structure, but it doesn’t have clearly demarcated screening, treatment and follow-up. Further, it only specifies specific lab tests at the start of the diagnosis and completion of therapy and lacks treatment period information. Often this table would require protocol digitization specialists to look at additional context (like surrounding texts on the page) in the protocol to determine whether or not it is a SoE table. The last table specifies pharmacokinetic collections and is not a SoE table (see prompts in appendix B which we wrote in consultation with the digitizers for SoE tables)